Vin Diesel owns the Fredonia Barber Shop. He employs 7 barbers and pays each a base rate of $1,640 per month. One of the barbers serves as the manager and receives an extra $510 per month. In addition to the base rate, each barber also receives a commission of $5.81 per haircut.

Other costs are as follows.
Advertising $200 per month
Rent $960 per month
Barber supplies $0.31 per haircut
Utilities $185 per month plus $0.30 per haircut
Magazines $30 per month

Vin currently charges $11.42 per haircut.
Determine the variable costs per haircut and the total monthly fixed costs. (Round variable costs to 2 decimal places, e.g. 2.25.)
Total variable cost per haircut $

Total fixed $

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

Compute the break-even point in units and dollars. (Round answers to 0 decimal places, e.g. 1,250.)
Break-even point

haircuts
Break-even point $

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

Determine net income, assuming 2,380 haircuts are given in a month.
Net income / (Loss) $

Click if you would like to Show Work for this question:

Open Show Work

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT
Answer #1
Per haircut
Barber Commission 5.81
Barber supplies 0.31
Utilities 0.30
Total variable cost per haircut 6.42
Cost Per month
Barber's salary ( 1640 * 7 ) 11480
Manager's salary 510
Advertising 200
Rent 960
Utilities 185
Magazines 30
Total fixed 13365
Contribution margin per unit = Service charge per haircut - Total variable cost per haircut = 11.42 - 6.42 5
Break-even point in units = Total fixed cost / Contribution margin per unit = 13365 / 5 2673 Haircuts
Break-even point in dollars = Break-even point in units * Service charges per haircut = 2673 * 11.42 30526
2380 haircuts
Service revenue ( 2380 * 11.42 ) 27180
(-) Variable cost ( 2380 * 6.42 ) 15280
Contribution margin 11900
(-) Fixed cost 13365
Net income / (loss) (1465)